LoopHOST

All Systems Operational

Data Center TQR-01 Operational
90 days ago
99.99 % uptime
Today
Links e Interconexões ? Operational
90 days ago
99.99 % uptime
Today
Equipamentos de Rede Operational
90 days ago
99.99 % uptime
Today
Estrutura Cloud Operational
90 days ago
100.0 % uptime
Today
Servidores de Armazenamento Operational
90 days ago
100.0 % uptime
Today
Servidores de Virtualização Operational
90 days ago
100.0 % uptime
Today
Servidores cPanel Operational
90 days ago
100.0 % uptime
Today
Servidores de Revenda ? Operational
90 days ago
100.0 % uptime
Today
Servidores de DNS ? Operational
90 days ago
100.0 % uptime
Today
Serviço de SMTP (Servidores cPanel) ? Operational
90 days ago
100.0 % uptime
Today
Estrutura VPS Operational
90 days ago
100.0 % uptime
Today
Servidores de Virtualização Operational
90 days ago
100.0 % uptime
Today
Plataformas de Atendimento Operational
90 days ago
100.0 % uptime
Today
SITE & ÁREA DO CLIENTE Operational
90 days ago
100.0 % uptime
Today
Área de Suporte Operational
90 days ago
100.0 % uptime
Today
Data Center POA2 Operational
90 days ago
100.0 % uptime
Today
Links e Interconexões Operational
90 days ago
100.0 % uptime
Today
Equipamentos de Rede Operational
90 days ago
100.0 % uptime
Today
Data Center SPO1 Operational
90 days ago
100.0 % uptime
Today
Links e Interconexões Operational
90 days ago
100.0 % uptime
Today
Equipamentos de Rede Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Aug 30, 2025

No incidents reported today.

Aug 29, 2025
Resolved - This incident has been resolved.
Aug 29, 01:59 GMT-03:00
Update - O problema de conectividade foi causado por uma falha de cordão optico MPO nos trunks de agregação de 200 G entre os roteadores de borda do data center e o switch CORE do data center. O switch CORE é responsável por efetuar a distribuição da conectividade para o restante do data center. O cordão com problema foi substituido e o acesso está normalizando.
Aug 29, 01:26 GMT-03:00
Identified - Identificamos uma falha de conectividade ao data center TQR-01 e TQR-02. Nossa equipe está investigando o problema.
Aug 29, 00:45 GMT-03:00
Aug 28, 2025

No incidents reported.

Aug 27, 2025

No incidents reported.

Aug 26, 2025

No incidents reported.

Aug 25, 2025

No incidents reported.

Aug 24, 2025

No incidents reported.

Aug 23, 2025

No incidents reported.

Aug 22, 2025

No incidents reported.

Aug 21, 2025

No incidents reported.

Aug 20, 2025

No incidents reported.

Aug 19, 2025

No incidents reported.

Aug 18, 2025

No incidents reported.

Aug 17, 2025

No incidents reported.

Aug 16, 2025

No incidents reported.